### v2

# Best Arm Identification for Prompt Learning under a Limited Budget

Abstract
The remarkable instruction-following capability of large language models (LLMs) has sparked a growing interest in automatically learning suitable prompts. However, while many effective methods have been proposed, the cost incurred during the learning process (e.g., accessing LLM and evaluating the responses) has not been considered. To overcome this limitation, this work explicitly incorporates a finite budget constraint into prompt learning. Towards developing principled solutions, a novel connection is established between prompt learning and fixed-budget best arm identification (BAI-FB) in multi-armed bandits (MAB). Based on this connection, a general framework TRIPLE (besT aRm Identification for Prompt LEarning) is proposed to harness the power of BAI-FB in prompt learning systematically. Unique characteristics of prompt learning further lead to two embedding-based enhancements of TRIPLE by exploiting the ideas of clustering and function approximation. Extensive experiments on multiple well-adopted tasks using both GPT 3.5 and Llama2 demonstrate the significant performance improvement of TRIPLE over the previous baselines while satisfying the limited budget constraints. Code available at https://github.com/ShenGroup/TRIPLE
1 Introduction
Large language models (LLMs) have rapidly changed technology landscapes in our society (Devlin et al.,, 2018; Touvron et al., 2023a, ; Touvron et al., 2023b, ; Bubeck et al.,, 2023; OpenAI, 2023a, ). Researchers continuously find effective ways to unlock their potential on various downstream tasks. Among different research directions, the remarkable ability of LLMs to follow instructions has motivated the study of searching for suitable prompts to interact with them (Liu et al.,, 2023). This approach is particularly attractive as it does not require updating the inside parameters of an LLM, and is natural in the way of human conversations. One representative example is the success of zero-shot Chain-of-Thoughts prompting (Kojima et al.,, 2022), i.e., adding a seemingly simple phrase of “let’s think step by step” bumps the reasoning capability of LLMs. Nevertheless, it has also been recognized that the performance of an LLM is sensitive to the selected prompts (Zhao et al.,, 2021; Lu et al.,, 2021), and manually designing suitable prompts can be a labor-intensive process (Mishra et al.,, 2021). Thus, there is a growing interest to automatically perform prompt learning, e.g., Zhou et al., (2022); Xu et al., (2022); Diao et al., (2022); Deng et al., (2022); Prasad et al., (2022); Zhang et al., 2023a ; Guo et al., (2023); Pan et al., (2023); Pryzant et al., (2023); Yang et al., (2023).
While these studies on prompt learning have gradually moved towards more practical scenarios of finding discrete and interpretable prompts with only black-box LLM accesses, the costs incurred during learning have not been considered. In particular, most of the previous works do not measure the required number of accesses to LLMs. However, in practice, accesses to LLMs are often (1) financially costly (e.g., each OpenAI API access incurs a cost); (2) time-wise consuming (e.g., even a locally hosted LLM would typically require seconds to respond); (3) under total usage limits (e.g., OpenAI has hard per-day and per-month limits on API accesses). Furthermore, it is often overlooked that evaluating the responses of an LLM for different prompts can be costly as many prompting tasks (e.g., writing improvement, mathematical reasoning, etc) would require human (and sometimes domain expert) opinions.
To address these challenges, this work proposes to study prompt learning under an explicitly imposed budget constraint when interacting with the targeted LLM, in addition to the previously considered constraints (e.g., discrete, interpretable, and black-box). To the best of our knowledge, budget constraints are only briefly mentioned in Zhou et al., (2022); Pryzant et al., (2023), and there are no systematic or principled investigations of how to address the limited budget constraint in prompt learning. The main contributions of this work are summarized as follows.
$\bullet$ The constraint of a limited budget is explicitly introduced into prompt learning, which has been largely ignored before. As most of the existing methods rely on selecting from a pre-generated candidate prompt pool, we focus our study on how to carefully allocate budgets to test each candidate prompt so that the optimal one can be learned efficiently and effectively.
$\bullet$ We develop a general solution framework, termed TRIPLE (besT aRm Identification for Prompt LEarning), by establishing a novel connection between prompt learning and multi-armed bandits (MAB) (Lattimore and Szepesvári,, 2020). In particular, we focus on harnessing the power of fixed-budget best arm identification (BAI-FB) (Audibert et al.,, 2010; Karnin et al.,, 2013) to address prompt learning (especially, selection) with a limited budget constraint. Two representative designs TRIPLE-SH and TRIPLE-CR, inspired by celebrated BAI-FB algorithms, are presented.
$\bullet$ To better handle large prompt pools and improve the scalability of our design, two enhanced methods, TRIPLE-CLST and TRIPLE-GSE, are proposed, where prompt embeddings are leveraged to accelerate learning. TRIPLE-CLST clusters prompts based on their embeddings for a two-phased scheme to first find the optimal cluster. TRILE-GSE, on the other hand, incorporates the idea of function approximation, i.e., sharing information among prompts via a learned function over their embeddings.
$\bullet$ Extensive experimental results are reported using well-adopted prompt tasks to demonstrate the superiority of the TRIPLE solutions over previous baselines. In particular, when handling small prompt pools, the basic TRIPLE-SH and TRIPLE-CR already outperform all prior baselines by (on average) $13\%$ to $35\%$; when tackling large prompt pools, the enhanced TRIPLE-CLST and TRIPLE-GSE further achieve additional $13\%$ to $16\%$ average gain over TRIPLE-CR, all on GPT 3.5. The gains are even more pronounced on Llama2-7b. Moreover, the proposed methods can be directly plugged into two popular prompt learning pipelines, APE (Zhou et al.,, 2022) and APO (Pryzant et al.,, 2023), with end-to-end performances significantly improved over their original implementations.
Besides the aforementioned technical contributions, we believe the value of this work extends broadly to providing a new perspective of prompt learning from MAB, and also a new application scenario of MAB in prompt learning. This established connection may spark further innovations in both fields as discussed in Sec. 7.
2 Prompt Learning under a Limited Budget
Following Zhou et al., (2022); Chen et al., (2023), we present a concrete formulation of the problem of prompt learning.
Consider that we are using an LLM $f(\cdot)$, which provides a mapping from any input $X\in{\mathcal{V}}$ to a distribution $\Delta_{{\mathcal{V}}}$ over the language space ${\mathcal{V}}$. The answer $\hat{Y}\in{\mathcal{V}}$ given by the LLM is assumed to be sampled from $f(X)$ as $\hat{Y}\sim f(X)$. Note that instead of treating $f(\cdot)$ as a deterministic function providing a specific output answer, we generally consider the practical setting where the answers of LLM exhibit a certain level of randomness.
For prompt learning, we aim to find a prompt $p$ such that when concatenated with inputs $X$ of a certain task (i.e, as $[p;X]$), it provides good performance in expectation with respect to the input distribution ${\mathcal{I}}_{X}$ and the inherent randomness of LLM $f(\cdot)$. The performance is measured as
$\mu(p):=\mathbb{E}_{X\sim{\mathcal{I}}_{X}}\mathbb{E}_{\hat{Y}\sim f([p;X])}\left[s(X,\hat{Y})\right],$
where $s(X,\hat{Y})$ denotes a score function that measures the quality of the output $\hat{Y}$ for the input $X$.
Motivated by the common usage scenario of LLMs, recent studies have imposed several constraints on this learning problem (Zhou et al.,, 2022; Pan et al.,, 2023; Chen et al.,, 2023; Guo et al.,, 2023), where the three key ones are
  • (I)
    Black-box: the method can be applied to black-box LLMs, i.e., only have access to an API $f(\cdot)$ and no access to the intermediate structure or parameters inside (including gradients, output likelihood, etc);
  • (II)
    Discrete: the learned prompt must be discrete characters, instead of continuous values (i.e., soft prompts);
  • (III)
    Interpretable: the learned prompt must be understandable by humans, instead of gibberish words.
Intuitively, the process of learning a good prompt would require interactions with the LLM (i.e., sample $\hat{Y}\sim f([p;X])$ and evaluating its responses (i.e., obtain score $s(X,\hat{Y})$). However, as mentioned in Sec. 1, such interactions and evaluations are often costly. Thus, besides the aforementioned constraints, we further explicitly take into account that the prompt learning process should have a limited budget:
  • (IV)
    Limited budget: the total number of trials with the LLM that happen during the learning is at most $N$.
Finally, the prompt learning problem considered in this work can be formulated as:
Finding $p^{*}$ with high performance $\mu(p^{*})$ under constraints of black-box, discrete, interpretable, and a limited budget.
Directly tackling this prompt learning problem has been widely recognized as challenging even without the constraint of a limited budget (Liu et al.,, 2023). As highlighted in Pryzant et al., (2023); Chen et al., (2023), it essentially requires performing a black-box discrete optimization. Instead, many proposed methods rely on the pipeline of first generating a pool of candidate prompts and then selecting from it (Jiang et al.,, 2020; Zhou et al.,, 2022; Xu et al.,, 2022; Prasad et al.,, 2022; Guo et al.,, 2023). The prompt generation can either be performed manually or follow designed automatic protocols. For example, the famous APE design (Zhou et al.,, 2022) selects from prompts generated by an LLM using demonstrations. From a unified perspective, we can simplify the problem into generating a pool of prompts ${\mathcal{P}}$ and finding the optimal prompt in it:
$p^{*}:=\operatorname*{arg\,max}_{p\in{\mathcal{P}}}\mu(p).$
While many efforts have been devoted along this line, we recognize that these previous methods (which are thoroughly reviewed in Sec. 6) are largely focused on how to generate prompts, while limited attention has been paid to how to select from the already generated prompts. For example, Jiang et al., (2020); Xu et al., (2022); Guo et al., (2023); Prasad et al., (2022) directly evaluate all the generated prompts on the entire development dataset; Zhou et al., (2022) adopts a naive filtering strategy without theoretical or empirical justification. Such treatment is understandable in previous works where budget limitations are not considered, i.e., unlimited evaluations can be performed for selection. With an explicit budget limitation, however, we need to carefully allocate the budgets to each prompt so that the optimal prompt (or at least a sufficiently good one) can be correctly learned, which is the main focus of this work. An overview of the considered prompt learning pipeline and our focus is illustrated in Fig. 1.
Figure 1: An overview of the commonly adopted prompt learning pipeline, which consists of two components: prompt generation and selection. Previous works mostly investigate the prompt generation component and ignore costs incurred during selection, where GrIPS and APE are proposed in Prasad et al., (2022); Zhou et al., (2022). This work, instead, mainly focuses on the prompt selection component, especially with an explicit budget constraint.
3 Connecting Prompt Learning with Best Arm Identification
We provide a new perspective of prompt learning through the lens of tools in multi-armed bandits (MAB) (Lattimore and Szepesvári,, 2020; Lai and Robbins,, 1985). In particular, prompt learning under a limited budget is shown to be intrinsically aligned with the problem of fixed-budget best-arm identification (BAI-FB) (Audibert et al.,, 2010; Karnin et al.,, 2013). In the following, a brief introduction to MAB is first provided. Then the connection between prompt learning (especially selection) with MAB (especially BAI-FB) is established. Based on this connection, we propose to both leverage existing and develop new BAI-FB designs to tackle the problem of prompt learning with a limited budget.
3.1 Multi-armed Bandits
The research of multi-armed bandits (MAB) has a long and rich history; see representative surveys of Lattimore and Szepesvári, (2020); Bubeck et al., (2012). The most basic form of MAB, i.e., the finite-armed stochastic bandits, considers a system with a set ${\mathcal{K}}$ finite arms (i.e., actions) that provide stochastic rewards when pulled. When interacting with the system, the agent can select one arm $k\in{\mathcal{K}}$ to pull at each time, and she receives a stochastic reward:
$r_{k}\sim\texttt{dist}_{k}(\nu_{k}),$
where $\texttt{dist}_{k}(\nu_{k})$ denotes the action $k$’s reward distribution with an unknown expectation $\nu_{k}$.
The learning objective of the agent in MAB can be roughly divided into two categories: (1) regret minimization, which maximizes the expected cumulative rewards collected by the agent (Auer et al.,, 2002; Audibert et al.,, 2009; Garivier and Cappé,, 2011); (2) best arm identification, which targets at outputting the best arm $k^{*}=\operatorname*{arg\,max}_{k\in{\mathcal{K}}}\nu_{k}$ (Audibert et al.,, 2010; Garivier and Kaufmann,, 2016; Jamieson and Nowak,, 2014). These two objectives often require different learning strategies. Regret minimization typically relies on a carefully designed balance between exploration (i.e., obtaining new information) and exploitation (i.e., collecting higher rewards based on the previous information). Best arm identification, on the other hand, is also called pure exploration as it only focuses on obtaining information to find the best arm.
3.2 A Bandit View of Prompt Learning
Based on the above introduction, it can be intuitively understood that the prompt learning (especially, selection) problem can be mapped into an MAB setting:
  • The pool of candidate prompts ${\mathcal{P}}$ is equivalent to the set of arms ${\mathcal{K}}$;
  • Using a prompt $p$ to interact with an LLM can be viewed as selecting a bandit arm $k$ to pull in the MAB model;
  • The feedback of the score function, i.e., $s(X,\hat{Y})$, provides the reward signal $r_{k}$, where $\texttt{dist}_{k}(\nu_{k})$ characterizes the randomness of $X\sim{\mathcal{I}}_{X}$ and $\hat{Y}\sim f([p;X])$. The expected performance $\mu(p)$ is the counterpart of the expected reward $\nu_{k}$ in MAB.
It can be further recognized that the target of prompt learning is more suitable to be captured as the best arm identification (BAI) problem, instead of a regret minimization one, as it only cares about finding the optimal prompt $p^{*}$ instead of the interaction performed during the learning process.
With the relationship between prompt learning and BAI established, we further consider the constraint of learning under a limited budget. We argue that this aligns with one of the main research directions in BAI called fixed-budget best arm identification (BAI-FB) (Karnin et al.,, 2013; Wang et al.,, 2023; Gabillon et al.,, 2012). BAI-FB particularly considers the problem of
maximizing the probability of correctly identifying the best arm $k^{*}$ while not pulling arms more than $T$ times.
It can be observed that this formulation matches the goal of prompt learning under a limited budget; thus BAI-FB provides us with a perfect toolbox to enhance the commonly required prompt selection process. The connection between prompt learning and MAB, in particular, BAI-FB, is further illustrated in Table 1. To avoid confusion, in the remainder of this paper, we will adopt the notation of prompt learning as introduced in Sec. 2.
Table 1: Connections between Prompt Learning and MAB.
Prompt Learning Multi-armed Bandits
The pool of prompts ${\mathcal{P}}$ The arm set ${\mathcal{K}}$
Interact LLM via prompt $p$ Pull arm $k$
Score $s(X,\hat{Y})$ Reward $r_{k}$
Randomness in $X$ and $\hat{Y}$ Randomness in $\texttt{dist}_{k}$
Performance $\mu(p)$ Expected reward $\nu_{k}$
Learn the optimal prompt under a limited budget Fixed-budget best arm identification (BAI-FB)
3.3 Harnessing the Power of BAI-FB
As mentioned, we recognize that prompt learning under a limited budget is a matching application scenario for BAI-FB. In this paper, we propose a general framework called TRIPLE (besT aRm Identification for Prompt LEarning) to harness the power of BAI-FB in solving the prompt learning problem. This is possible because BAI-FB has witnessed significant development over the years, with several efficient designs being proposed. As a first step, we choose two popular and successful BAI-FB schemes and implement them for prompt learning, which are briefly described below.
Sequential Halving (SH) (Karnin et al.,, 2013): SH is one of the first provably efficient BAI-FB designs and remains popular after a decade of its proposal. It follows a protocol that divides the total budget $N$ into $\lceil\log_{2}(|{\mathcal{P}}|)\rceil$ equal-length phases. In each phase, SH uniformly tries all active prompts (initialized as ${\mathcal{P}}$) and eliminates half of them with the lower sample means for the next phase. The final active arm is output as the identified optimal prompt. A complete description of TRIPLE-SH is provided in Alg. 2 of Appendix A.
Continuously Reject (CR) (Wang et al.,, 2023): CR is a recently proposed method, which can be viewed as an extension of the classical Successively Reject (SR) design (Audibert et al.,, 2010). It uniformly explores active prompts (initialized as ${\mathcal{P}}$) and performs potential elimination of poorly-performed prompts after each pull. The elimination is based on carefully designed criteria using the Large Deviation Principle. It can be observed that, without the phased structure, CR is more adaptive than SH (and SR), which makes it appealing both theoretically and practically. We present details of TRIPLE-CR in Alg. 3 of Appendix A.
While MAB has found broad applications in recommender systems (Li et al.,, 2010), healthcare (Shen et al.,, 2020), wireless communications (Gai et al.,, 2012), and beyond (Bouneffouf and Rish,, 2019), a systematical connection between MAB and prompt learning has not been established before to the best of our knowledge, which may spark new research activities (see discussions in Sec. 7). In addition, although SH and CR are selected as the representatives, the connection between prompt learning and MAB is fundamental, e.g., any existing or forthcoming BAI-FB designs can be flexibly incorporated into TRIPLE.
4 Handling Large Candidate Pools via Prompt Embeddings
The connection established in the last section provides us with the core idea of leveraging BAI-FB designs to tackle prompt learning. As having been theoretically established (Audibert et al.,, 2010), solving BAI-FB without additional structures, however, will unavoidably incur an identification error that is positively related to the number of candidate prompts $|{\mathcal{P}}|$. In other words, given a larger pool of prompts, it becomes harder to find the optimal prompt with the basic BAI-FB designs, which restricts their applicability to practical prompt learning problems.
The key reason behind this is that each candidate prompt is treated independently in the basic BAI-FB. Thus, budgets need to be assigned to all of the prompts and no information can be shared among them, which is often not the case in prompt learning. For a prompt learning problem, the underlying task is often stable, e.g., rewriting emails, constructing TLDR, etc. The candidate prompts, regardless of their generation methods, should all reflect the purpose of the underlying task and thus share similarities. For example, the candidate prompts generated via demonstrating LLMs (Zhou et al.,, 2022) often share similar structures and differ only in a few words or word orders.
With the above observation, we target sharing information among prompts during learning. To achieve this, we propose to leverage an embedding model, denoted as $\texttt{embed}:{\mathcal{V}}\to{\mathbb{R}}^{d}$, to obtain the sentence embedding of the prompts:
$e(p):=\texttt{embed}(p)\in{\mathbb{R}}^{d},~{}~{}{\mathcal{E}}:=\{e(p):p\in{\mathcal{P}}\},$
where $d$ refers to the embedding dimension. In the experiments, the OpenAI embedding API is adopted for embedding while, in general, any sufficiently expressive models can be incorporated. Also, due to this flexibility, using embedding models is fundamentally different from requiring a white-box LLM as in Chen et al., (2023); Lin et al., (2023).
With the obtained prompt embeddings, we propose two useful enhancements to further improve the learning effectiveness when the pool of candidate prompts is large.
Algorithm 1 TRIPLE-CLST
1:  Input: the pool of candidate prompts ${\mathcal{P}}$ and their embeddings ${\mathcal{E}}$, overall budget $N$, Phase I budget $N_{1}$, number of clusters $L$
2:  Cluster ${\mathcal{P}}$ into clusters ${\mathcal{C}}=\{{\mathcal{C}}^{1},\cdots,{\mathcal{C}}^{L}\}$ based on embeddings ${\mathcal{E}}$ (e.g., via $k$-means)
3:  Obtain $\widehat{{\mathcal{C}}}^{*}\leftarrow\text{BAI-FB}({\mathcal{C}},N_{1})$Phase I}
4:  Obtain $\hat{p}^{*}\leftarrow\text{BAI-FB}(\widehat{{\mathcal{C}}}^{*},N-N_{1})${Phase II}
5:  Output: prompt $\hat{p}^{*}$
4.1 Leveraging Similarities via Clustering
Since the key challenge is a large pool of candidate prompts, an intuitive idea is to effectively decrease the size of the pool. We thus propose a two-phased BAI-FB scheme for prompt learning. In Phase I, the entire pool of candidate prompts is clustered into several groups based on their embeddings, and BAI-FB is performed on the clusters with an initial target of finding the optimal cluster (or the few good clusters). Then, in Phase II, BAI-FB is performed on the prompts in the optimal cluster with the target of identifying one final prompt. For both phases, different BAI-FB designs can be incorporated, e.g., SH and CR. The entire procedure, referred to as TRIPLE-CLST, is described in Alg. 1.
The effectiveness of TRIPLE-CLST relies on the clustering results produced in Phase I. Ideally, prompts with similar performances should be clustered together. In this way, Phase I can quickly eliminate the prompts with poor performances, leaving a small pool of good prompts for Phase II to process. In the experiments, this intuitive phenomenon is indeed observed. Here we provide examples in Fig. LABEL:fig:cluster (with more details in Appendix C.1), where as expected, prompts in the same cluster share similar performances.
4.2 Sharing Information via Function Approximation
Besides clustering, another idea to incorporate the prompt features is to learn a common function (e.g., an MLP) to predict the prompt features based on their features. Similar ideas of function approximation have also been widely adopted in MAB literature to share information among large action spaces, with function classes ranging from linear functions (Abbasi-Yadkori et al.,, 2011; Yang and Tan,, 2022) to neural networks (Zhu et al.,, 2021; Zhou et al.,, 2020).
In the setting considered in this work, we adopt a recently developed BAI-FB scheme as described in the following:
GSE (Azizi et al.,, 2023): It inherits the general phased elimination flow of SH described in 3.3. The major difference is that SH uses sample means to perform eliminations. GSE, on the other hand, leverages collected samples from previous phases to train a reward function $g_{\bm{\theta}}(\cdot):{\mathbb{R}}^{d}\to{\mathbb{R}}$ that maps prompt embeddings to the predicted performance, which is further used to eliminate prompts. The details of TRIPLE-GSE are provided in Alg. 4 of Appendix A.
5 Experiments
In this section, extensive experimental results are reported to evaluate the efficiency of TRIPLE across diverse prompting tasks from two standard datasets: Instruction-Induction (Honovich et al.,, 2022) and BigBench (Suzgun et al.,, 2022). The results reported in this section are collected from gpt-3.5-turbo-1106 model of OpenAI (OpenAI, 2023a, ) as well as Llama2-7b (Touvron et al., 2023b, ). Full experimental details can be found in Appendix B. We particularly focus on $12$ representative tasks, while the complete results of the overall $47$ tasks are reported in Appendix C.
5.1 Evaluating TRIPLE with Fixed Prompt Pools
As TRIPLE main focuses on the prompt selection component, we perform initial evaluations in an isolated fashion of selecting from fixed pools of candidate prompts. For this experiment, candidate pools of prompts are generated following the well-established APE design (Zhou et al.,, 2022) with a high LLM temperature to ensure randomness. Then, under a limited budget, the performances of TRIPLE algorithms are compared with the following two baselines:
  • Uniform. Many previous designs choose to evaluate the entire candidate pool on all development data, e.g., Guo et al., (2023); Prasad et al., (2022). With a limited budget, this strategy corresponds to a uniform allocation, i.e., equally dividing the total budget to test all prompts.
  • UCB. The upper confidence bound (UCB) method is a famous design for regret minimization in MAB. Its variants have been applied to prompt selection in Pryzant et al., (2023); Lin et al., (2023). We evaluate UCB using its vanilla version from Auer et al., (2002).
Performance with small prompt pools. We begin with relatively small candidate pools, generating $10$ prompts for selection in each task. Results with an overall budget of $50$ are reported in Fig. LABEL:fig:bar_fig_main (top). It can be observed that TRIPLE methods outperform two baselines on most of the tasks, evidencing its prowess in allocating limited budgets for prompt selection. Especially, across these tasks, TRIPLE-CR consistently outperforms Uniform with improvements ranging from $8\%$ to $129\%$, and on average, a notable gain of $35\%$. Moreover, it achieves a $3\%$ to $30\%$ improvement over UCB, with an average $13\%$ gain. The benefit of using TRIPLE over Uniform is intuitive as BAI-FB designs are much more adaptive. The improvement over UCB further highlights the ineffectiveness of methods for regret minimization in prompt learning and provides empirical justifications for using BAI-FB.
Performance with large prompt pools. We next test larger candidate pools with $30$ prompts per task. Results reported in Fig. LABEL:fig:bar_fig_main (middle) reflect the selection performance with an overall budget of $150$. It can be observed that while TRIPLE-SH and TRIPLE-CR still achieve better performance than Uniform and UCB, the enhanced TRIPLE-CLST and TRIPLE-GSE demonstrate more remarkable improvements, evidencing their capabilities in leveraging prompt embeddings to handle large prompt pools. In particular, both TRIPLE-CLST and TRIPLE-GSE significantly outperform the previously best TRIPLE-CR method (let alone Uniform and UCB), with a gain ranging from $2\%$ to $60\%$, averaging of $13\%$ and $16\%$, respectively. With the same prompt pools, we also test the performance on Llama2-7b with results shown in Fig. LABEL:fig:bar_fig_main (bottom). We can see that the superiority of TRIPLE is even more evident on Llama-2: TRIPLE-GSE and TRIPLE-CLST achieve average gains of $45\%$ and $51\%$ over Uniform across all tasks, reaching a remarkable $200\%$ improvement on the “larger__\__animal” task.
Impact of the total budget. For a more comprehensive understanding, using candidate pools with $30$ prompts, we further examine the impact of budgets, starting with $5$ evaluations per prompt on average (i.e., $150$ overall as adopted in Fig. LABEL:fig:bar_fig_main (middle)), and then gradually increasing to $30$ (i.e., $900$ overall, which is the same as the experiments in Zhou et al., (2022)). From the results shown in Fig. LABEL:fig:budget (left), we see that the improvements of TRIPLE over baselines are more pronounced with lower budgets. In particular, with a budget of $10$ evaluations per prompt on average, TRIPLE-CR, TRIPLE-CLST and TRIPLE-GSE maintain notable $9.7\%$, $13.5\%$ and $17.4\%$ improvement over Uniform, respectively; when the budget escalates to $20$ evaluations per prompt on average, TRIPLE-CLST and TRIPLE-GSE still achieve an approximate $8\%$ improvement. Once the budget reaches $30$ evaluations per prompt on average, all methods provide approximately the same performance as they can all identify the optimal prompts under this generous budget.
To further guide practical implementation, we additionally investigate how to select a reasonable budget. In particular, we focus on the efficiency of various prompt selection algorithms in identifying a “good” prompt – either the optimal prompt in the pool or achieving $95\%$ or better of the optimal prompt’s performance. Fig. LABEL:fig:budget (right) illustrates that initial increases in budgets significantly improve the probability of identifying a good prompt, but this benefit tapers off with further budget expansion. This finding suggests that starting with a modest budget and incrementally increasing it is the more effective approach, stopping when additional investment no longer translates into significant returns.
Table 2: Performances of integrating TRIPLE in the end-to-end pipelines. The baseline methods reported in the original implementations are labeled as (b). For each task, the best score across two pipelines is marked as red, and the best score in the remaining pipeline is highlighted as yellow. TRIPLE-CR are selected over TRIPLE-SH due to its better performance observed in the previous experiments. TRILE-CLST is ignored in the tests with APO, as it is ineffective to cluster only $10$ prompts.
APE (Zhou et al.,, 2022) APO (Pryzant et al.,, 2023)
Tasks Uniform (b) CR CLST GSE UCB (b) CR GSE
Cause and effect 0.65 $\pm$ 0.18 0.74 $\pm$ 0.06 0.75 $\pm$ 0.13 0.78 $\pm$ 0.08 0.78 $\pm$ 0.15 0.80 $\pm$ 0.05 0.80 $\pm$ 0.08
Common concept 0.09 $\pm$ 0.05 0.12 $\pm$ 0.06 0.10 $\pm$ 0.04 0.14 $\pm$ 0.05 0.12 $\pm$ 0.04 0.12 $\pm$ 0.05 0.14 $\pm$ 0.01
Disambiguation qa 0.83 $\pm$ 0.04 0.88 $\pm$ 0.10 0.97 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.04 0.98 $\pm$ 0.02 0.96 $\pm$ 0.02
Gender inc. DE 0.74 $\pm$ 0.17 0.81 $\pm$ 0.10 0.85 $\pm$ 0.12 0.84 $\pm$ 0.14 0.69 $\pm$ 0.22 0.80 $\pm$ 0.17 0.88 $\pm$ 0.05
Hyperbaton 0.78 $\pm$ 0.07 0.83 $\pm$ 0.11 0.84 $\pm$ 0.12 0.84 $\pm$ 0.11 0.59 $\pm$ 0.24 0.74 $\pm$ 0.21 0.79 $\pm$ 0.18
Larger animal 0.56 $\pm$ 0.24 0.64 $\pm$ 0.25 0.79 $\pm$ 0.06 0.84 $\pm$ 0.02 0.66 $\pm$ 0.13 0.73 $\pm$ 0.18 0.85 $\pm$ 0.15
Movie recommendation 0.61 $\pm$ 0.12 0.65 $\pm$ 0.18 0.76 $\pm$ 0.06 0.74 $\pm$ 0.14 0.67 $\pm$ 0.11 0.65 $\pm$ 0.15 0.71 $\pm$ 0.15
Object counting 0.41 $\pm$ 0.12 0.45 $\pm$ 0.08 0.50 $\pm$ 0.07 0.48 $\pm$ 0.12 0.44 $\pm$ 0.08 0.50 $\pm$ 0.09 0.49 $\pm$ 0.07
Orthography starts with 0.41 $\pm$ 0.21 0.65 $\pm$ 0.16 0.67 $\pm$ 0.12 0.66 $\pm$ 0.13 0.58 $\pm$ 0.13 0.64 $\pm$ 0.09 0.67 $\pm$ 0.17
Question selection 0.90 $\pm$ 0.04 0.91 $\pm$ 0.03 0.95 $\pm$ 0.01 0.93 $\pm$ 0.03 0.93 $\pm$ 0.06 0.92 $\pm$ 0.06 0.93 $\pm$ 0.03
Rhymes 0.66 $\pm$ 0.30 0.68 $\pm$ 0.26 0.75 $\pm$ 0.20 0.78 $\pm$ 0.16 0.78 $\pm$ 0.12 0.83 $\pm$ 0.08 0.85 $\pm$ 0.13
Snarks 0.44 $\pm$ 0.10 0.52 $\pm$ 0.19 0.57 $\pm$ 0.10 0.60 $\pm$ 0.21 0.49 $\pm$ 0.17 0.56 $\pm$ 0.15 0.67 $\pm$ 0.05
5.2 Integrating TRIPLE into End-to-End Pipelines
We now explore whether TRIPLE can provide performance improvements when plugged into end-to-end prompt learning pipelines that include both prompt generation and selection. To this end, two end-to-end designs are considered, aiming to assess the performance of TRIPLE in more fluid and iterative settings, which are discussed in the following with our implementation details.
  • APE. Proposed by Zhou et al., (2022), the APE pipeline lets LLMs generate prompt candidates and then selects from them. In our experiments, for each task, following original templates, $30$ prompts are generated, followed by different methods to perform selection with a budget of $5$ evaluations per prompt on average (i.e., $150$ LLM accesses overall). The authors of Zhou et al., (2022) suggest a non-iterative version with uniform evaluations of prompts, which is taken as the baseline here.
  • APO. The APO pipeline (Pryzant et al.,, 2023) is an iterative one, letting LLMs criticize the previous prompts. Here, following the original templates, three iterations are performed and $10$ prompts are generated per iteration. Different selection methods are then tested with a budget of $50$ per iteration so that an overall budget of $150$ is used, aligning with that of APE. Pryzant et al., (2023) have reported UCB as the most effective prompt selection method, which is adopted as the baseline here.
The end-to-end experiment results are reported in Table 2, which reveal the consistently better performance of TRIPLE over the originally adopted baseline methods. This observation highlights the applicability and flexibility of the TRIPLE framework, i.e., it can benefit any prompt learning pipelines requiring a selection component.
6 Related Works
Prompt Learning. The study of prompt learning (also known as instruction learning) focuses on automatically learning suitable prompts, which is more scalable compared with manual prompt engineering. Many efforts have been devoted to this direction (Gao et al.,, 2020; Wen et al.,, 2023; Sun et al.,, 2023; Zhang et al.,, 2022; Zhang et al., 2023b, ), which is summarized in a recent survey of Liu et al., (2023). The early studies mostly consider optimizing soft prompts (i.e., continuous vectors) (Lester et al.,, 2021; Li and Liang,, 2021; Zhong et al.,, 2021; Liu et al.,, 2021; Wu et al.,, 2022) or discrete but not interpretable prompts (Shin et al.,, 2020; Shi et al.,, 2022) for white-box LLMs, where different gradient-guided optimization techniques are leveraged. The recent research efforts, instead, are more focused on considering the more practical setting of learning interpretable prompts for black-box LLMs (Prasad et al.,, 2022; Zhou et al.,, 2022; Pryzant et al.,, 2023; Xu et al.,, 2022; Guo et al.,, 2023; Pan et al.,, 2023; Chen et al.,, 2023), where the generating-then-selecting pipeline in Fig. 1 is often adopted.
Compared with previous investigations, this work additionally considers a limited budget, which we believe is a practical but under-investigated concern. Most of the previous methods perform selection based on evaluating prompts on all development data (Jiang et al.,, 2020; Xu et al.,, 2022; Guo et al.,, 2023; Prasad et al.,, 2022), which is unavoidably costly. APE (Zhou et al.,, 2022) briefly touched on the cost issue by proposing a naive iterative top filtering strategy (which however is suggested to have only marginal benefits). APO (Pryzant et al.,, 2023) tested a few MAB methods, including two classical BAI-FB designs, i.e., SH (Karnin et al.,, 2013) and SR (Audibert et al.,, 2009), and reported that the performance of UCB is the more favorable. However, it neither formally introduces the budget constraint nor provides a systematical connection to BAI-FB as in this work. Also, this work goes much deeper in incorporating the state-of-the-art BAI-FB designs (i.e., CR (Wang et al.,, 2023) and GSE (Azizi et al.,, 2023)) and proposing embedding-based enhancements. It is worth noting that INSTINCT (Lin et al.,, 2023) also incorporates one MAB method, i.e., NeuralUCB (Zhou et al.,, 2020), to prompt learning. However, Lin et al., (2023) requires a white-box LLM while NeuralUCB is designed for regret minimization (instead of best arm identification), which is not suitable for learning the optimal prompt.
Multi-armed Bandits. Here we briefly discuss the representative studies of MAB, with comprehensive surveys available in Bubeck et al., (2012); Lattimore and Szepesvári, (2020). As mentioned in Sec. 3.1, the target of learning in a MAB system can be roughly categorized as regret minimization and best arm identification. The regret minimization designs target achieving a desired balance between exploration and exploitation so that the cumulative rewards are maximized, e.g., UCB (Auer et al.,, 2002) and Thompson sampling (Thompson,, 1933). The best arm identification (BAI) designs are fully focused on exploration and can be further divided into two classes: fixed-budget (BAI-FB) and fixed-confidence (BAI-FC). The BAI-FB setting maximizes the probability of finding the best arm with a limited number of pulls (Audibert et al.,, 2010; Karnin et al.,, 2013; Wang et al.,, 2023; Azizi et al.,, 2023; Atsidakou et al.,, 2022; Yang and Tan,, 2022). The BAI-FC setting is a dual one which focuses on minimizing the overall number of pulls while guaranteeing that the probability of finding the best arm is higher than a threshold (Garivier and Kaufmann,, 2016; Jamieson and Nowak,, 2014; Soare et al.,, 2014; Jedra and Proutiere,, 2020). Besides leveraging BAI-FB as in this work, it is imaginable that BAI-FC designs can also find applications in prompt learning, especially when valuing identification accuracy over incurred costs. While MAB has found wide success in different applications, this work marks the first time that a systematical connection between MAB and prompt learning has been established to the best of our knowledge.
7 Conclusions
Prompt learning is an important problem for large language models (LLMs), but prior research has not considered the potential cost during prompt selection. We have explicitly incorporated a budget constraint to prompt learning, and studied the problem of how to efficiently select prompts with the given budget. A systematical connection between multi-armed bandits (MAB) and prompt learning was established. Through this lens, we proposed a general framework, termed TRIPLE, to fully harness the power of fixed-budget best arm identification (BAI-FB) to perform prompt learning. Besides standard BAI-FB designs, two embedding-based enhancements were proposed to accelerate learning when selecting from large prompt pools. Extensive experimental results demonstrated the superiority of TRIPLE on prompt selection in multiple representative tasks and various targeted LLMs. Furthermore, we showed that TRIPLE could be plugged into popular end-to-end prompt learning pipelines, with better performance than previous implementations, demonstrating its effectiveness and flexibility.
In addition to the technical contributions, we believe that the connection between prompt learning and MAB may be of broader interest. It not only provides a rich set of tools from MAB to advance prompt learning but also introduces a new application scenario for MAB (especially best arm identification) research. For example, the research on contextual bandits (Li et al.,, 2010) may provide insights into selecting input-dependent prompts (Wu et al.,, 2022). Also, the new application of prompt learning may spark new research efforts in MAB, e.g., developing more efficient BAI methods for correlated arms.
Appendix A Details of the Adopted BAI-FB Designs
The details of TRIPLE-SH (inspired by Karnin et al., (2013)), TRIPLE-CR (inspired by Wang et al., (2023)), and TRIPLE-GSE (inspired by Azizi et al., (2023)) can be found in Algs. 2, 3, and 4, respectively.
Algorithm 2 TRIPLE-SH
1:  Input: the pool of candidate prompts ${\mathcal{P}}$, budget $N$
2:  Initialization: set $\hat{\mu}(p)\leftarrow$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
3:  for phase $p=1,\cdots,\lceil\log_{2}(|{\mathcal{P}}|)\rceil$ do
4:     Interact with the targeted LLM using each prompt in ${\mathcal{A}}$ for $\left\lceil N/(|{\mathcal{A}}|\lceil\log_{2}(|{\mathcal{P}}|)\rceil)\right\rceil$ times
5:     Update the sample means $\{\hat{\mu}(p):p\in{\mathcal{A}}\}$ using the collected samples
6:     Update the active prompt set ${\mathcal{A}}$ as the set of $\lceil{\mathcal{A}}/2\rceil$ prompts in the original ${\mathcal{A}}$ with the highest $\hat{\mu}(p)$
7:  end for
8:  Output: the remaining active prompt $\hat{p}^{*}$
Algorithm 3 TRIPLE-CR
1:  Input: the pool of candidate prompts ${\mathcal{P}}$, budget $N$
2:  Initialization: set $n(p)\leftarrow 0$, $\hat{\mu}(p)\leftarrow$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
3:  for time $\tau=1,\cdots,N$ do
4:     Receive input $x_{\tau}$
5:     Select prompt $p_{\tau}\leftarrow\operatorname*{arg\,min}_{p\in{\mathcal{A}}}n(p)$
6:     Sample output $\hat{y}_{\tau}\sim f([p_{\tau},x_{\tau}])$ from the targeted LLM
7:     Obtain score $s_{\tau}\leftarrow s(x_{\tau},\hat{y}_{\tau})$
8:     Update $\hat{\mu}(p_{\tau})\leftarrow\frac{\hat{\mu}(p_{\tau})n(p_{\tau})+s_{\tau}}{n(p_{\tau})+1}$ and $n(p_{\tau})\leftarrow n(p_{\tau})+1$
9:     Compute $p^{\prime}\leftarrow\operatorname*{arg\,min}_{p\in{\mathcal{A}}}\hat{\mu}(p)$ and $\delta_{\tau}\leftarrow\frac{\sum_{p\in{\mathcal{A}}\backslash\{p^{\prime}\}}\hat{\mu}(p)}{|{\mathcal{A}}|-1}-\hat{\mu}(p^{\prime})$
10:     if $\sqrt{\frac{N-\sum_{p\notin{\mathcal{A}}}n(p)}{\sum_{p\in{\mathcal{A}}}n(p)\log(|{\mathcal{A}}|)}}-1\leq\delta_{\tau}$ then
11:        Eliminate prompt $p^{\prime}$, i.e., ${\mathcal{A}}\leftarrow{\mathcal{A}}\backslash\{p^{\prime}\}$
12:     end if
13:  end for
14:  Output: prompt $\hat{p}^{*}\leftarrow\operatorname*{arg\,max}_{p\in{\mathcal{A}}}\hat{\mu}(p)$
Algorithm 4 TRIPLE-GSE
1:  Input: the pool of candidate prompts ${\mathcal{P}}$ and their embeddings ${\mathcal{E}}$, budget $N$
2:  Initialization: set $\hat{\mu}(p)\leftarrow$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
3:  for phase $p=1,\cdots,\lceil\log_{2}(|{\mathcal{P}}|)\rceil$ do
4:     Interact with the targeted LLM using each prompt in ${\mathcal{A}}$ for $\left\lceil N/(|{\mathcal{A}}|\lceil\log_{2}(|{\mathcal{P}}|)\rceil)\right\rceil$ times
5:     Use the collected samples to train a function $g_{\bm{\theta}}(\cdot)$ parameterized by ${\bm{\theta}}$, e.g., a linear function or an MLP
6:     Compute $\{\hat{\mu}(p)=g_{\bm{\theta}}(e(p)):p\in{\mathcal{A}}\}$
7:     Update the active prompt set ${\mathcal{A}}$ as the set of $\lceil{\mathcal{A}}/2\rceil$ prompts in the original ${\mathcal{A}}$ with the highest $\hat{\mu}(p)$
8:  end for
9:  Output: the remaining active prompt $\hat{p}^{*}$
Appendix B Full Experimental Details
In this section, we present full details of our experiments. Prior to delving deeper into the discussion, as we use chat-based LLMs, initial system instructions are needed. We adopt the officially recommended system instructions in the experiments, as shown in Fig LABEL:fig:sys_instruction.
B.1 The Adopted Score Functions: Metrics for Evaluation
Different score functions $s(\cdot,\cdot)$, i.e., metrics for evaluation, are used for diverse tasks in the Instruction-Induction and BigBench-ii datasets, namely “Exact match”, “F1-score”, “Multiple choice within”, and “Multiple choice f1-score”. These score functions are adopted according to the specific output requirements of different tasks:
  • Exact match: Used for most tasks unless otherwise specified, this metric scores $1$ for outputs exactly matching the label, and 00 otherwise.
  • f1-score: Applied to tasks with complex targets like long sentences (e.g., “informal to formal”, “negation”, “sentence similarity”), this metric (defined in Definition B.1), evaluates the overlap between the LLM response and the label.
  • Multiple choice within: Suitable for tasks with several correct answers, it scores $1$ if the LLM’s response matches any correct answer and 00 otherwise. We utilized this metric for tasks “rhymes”, “translation-en-de”, “translation-en-es”, “translation-en-fr” and “word in context”.
  • Multiple choice f1-score: Employed for tasks with multiple, lengthy correct answers (“common concept” task), it calculates the highest f1-score across all potential correct answers.
Definition B.1 (f1-score).
Suppose the question has a labeled answer $T$ and the response of the LLM is $A$, then the f1-score for this answer is defined as:
$S_{f1}=\frac{2\times P\times R}{P+R},$
where $P=l_{m}/l_{A}$ stands for the precision of the response and $R=l_{m}/l_{T}$ the recall of the response. Here we use $l_{A}$ and $l_{T}$ to denote the length of the response and label while $l_{m}$ is adopted to represent the number of matching words between $A$ and $T$.
For the specific score function adopted for the BigBench-ii dataset, we advise referring to the “metric” label for each task therein. This label indicates the appropriate metric (“Exact match” or “Multiple choice within”) for the optimal evaluation.
B.2 Experiments with Fixed Prompt Pools and APE
The prompt generation process to obtain the fixed prompt pools largely follows the one in APE (Zhou et al.,, 2022), i.e., demonstrating LLMs with examples. In particular, in the generation of each prompt, we sample 10 examples from the training set to demonstrate LLMs with two types of generation templates: ‘forward’ and ‘backward’, which are illustrated in Fig. LABEL:fig:generation_temp. The same setups are also adopted in the end-to-end experiments with APE in Sec. 5.2. A side observation is that we find that in general, GPT-3.5 can handle both templates, resulting in reasonable prompts. However, Llama2-7b exhibits difficulties in generating useful prompts from the ‘backward’ template, possibly due to its more simplified structure.
B.3 Experiments with APO
[…]
B.4 Implementions of TRIPLE-CLST and TRIPLE-GSE
Obtaining Embedding. A critical component of both TRIPLE-CLST and TRIPLE-GSE is the extraction of sentence embeddings for the candidate prompts. In our experiments, the prompts are first tokenized using the cl100k__\__base tokenizer. Then, the tokenized prompts are input into the text-embedding-ada-002 model (OpenAI, 2023b, ), converting them into continuous vectors.
TRIPLE-CLST. In experiments with TRIPLE-CLST, the number of clusters is set as $L=\lceil\sqrt{|{\mathcal{P}}|}\rceil$ and a third of our total budget is allocated for the initial phase, i.e., $N_{1}=N/3$. The $k$-means algorithm is employed as the clustering method. For more stable performance, Phase I is configured to find the top $L/2$ clusters, instead of the optimal one, which safeguards against the situation that the optimal prompt is not located in the optimal cluster. Also, for the BAI-FB designs in both phases, the CR algorithm (Wang et al.,, 2023) is adopted due to its flexibility.
[…]
Appendix C Additional Experimental Results
In this section, additional experimental results are provided to supplement observations in the main paper.
C.1 Additional Details of Clustering
In Sec. 4, clustering is adopted to enhance the efficiency of prompt selection in TRIPLE-CLST. To supplement the two examples provided in Fig. LABEL:fig:cluster, Tables 3 and 4 provide the full details of the prompts to be clustered and showcase prompts that have been grouped into the same clusters. These examples reveal that clustered prompts often share thematic elements or keywords, such as the use of ‘choice/choose’ and ‘movie’ in the movie selection task, Cluster 0, or a focus on ‘rhyming words’ in the rhymes task, Cluster 3. In addition, when the prompt pools are small, we observe that the clustering process struggles to create balanced and meaningful clusters. Thus, the results of TRIPLE-CLS with small prompts pools with $|{\mathcal{P}}|=10$ are not reported. Addressing this challenge could involve developing methods to enforce balanced clustering or innovatively reallocating the budget for more efficient phase I cluster identification. These considerations open avenues for future research.
C.2 Impacts of Expanding Prompt Pools
One additional interesting phenomenon observed is that the expansion of the prompt pool can transform the nature of tasks, oscillating between easy and hard. Specifically, as depicted in Fig. LABEL:fig:change_pool (top), in most tasks, increasing the number of prompts often shifts tasks from easy to hard, i.e., with one clearly better prompt in the pool with size $10$ while having several comparably good prompts in the pool with size $30$. Conversely, as shown in Fig. LABEL:fig:change_pool (bottom), in certain tasks, expanding the prompt pool actually highlights the good prompts.
C.3 Complete Evaluations on $47$ Tasks
In the main paper, we provide experimental results on $12$ representative tasks from the overall $47$ available tasks. In the following, the complete results are discussed.
Instruction-Induction. Results on the $24$ available tasks in the Instruction-Induction dataset (Honovich et al.,, 2022) are illustrated in Fig. 2 (small prompt pools $|{\mathcal{P}}|=10$ with gpt-3.5), 3 (large prompt pools $|{\mathcal{P}}|=30$ with gpt-3.5), and 4 (large prompt pools $|{\mathcal{P}}|=30$ with llam2-7b). It can be observed that TRIPLE methods improve Uniform on most tasks while achieving matching performance on the remaining ones.
BigBench-ii. Results on the $23$ available tasks in the BigBench-ii dataset (Suzgun et al.,, 2022) are reported in Figs. 5 (small prompt pools $|{\mathcal{P}}|=10$ with gpt-3.5), 6 (large prompt pools $|{\mathcal{P}}|=30$ with gpt-3.5), and 7 (large prompt pools $|{\mathcal{P}}|=30$ with Llam2-7b). It can still be observed that TRIPLE methods consistently surpass or at least match the performance of baselines.
Besides the average return, another key aspect of prompt selection is the frequency of selecting a good prompt. In Table 5, we further demonstrate the best prompt identification frequency of different algorithms across $20$ selected tasks from 5 independent runs.
Table 3: Clusters for “movie selection”: the best prompt overall is marked in red, and the best prompt in each cluster in yellow.
Cluster
Prompts
0
The instruction was to select the movie title that appeared the most frequently among the given choices.
The instruction was to choose the correct movie from the given choices based on the input movie titles.
Based on the given inputs and choices, the task was to select the option that matched the given genre film.
The instruction was to choose the movie title from the given choices that corresponds to the movie titles in the input.
The instruction was to select the movie from the given choices.
The instruction was to determine which movie from a list of choices the user should watch based on the inputted movies.
The instruction was to select the movie title that received the most number of votes from the given choices.
The instruction was to choose the correct movie based on the given choices.
The instruction was to provide the output movie based on the given choices.
The instruction was to choose one movie from the given choices.
The instruction was to select the correct movie from the given list of choices.
The instruction was to provide the output movie choice for each given input movie titles and their corresponding choices.
The instruction was to recommend one movie out of the given choices.
The instruction given was to determine the best choice from a given list of movies, based on a set of choices and their corresponding scores.
The instruction was to choose the most suitable movie choice based on given input movies. There were multiple choices for each input, and the selected choice was the one with a value of 1. The chosen movie is the output.
1 Choose the correct answer based on the given choices.
Choose the correct answer from the given choices.
2 In each case, provide output responding the relative place of these Nos among Fraser beat shape singers then,the relative here is taken negatively (((( greater– worser place’s id it falls in franchise with counselling distribution)))) Crush Orders Miscellaneous similarly ) depending continuity concentration tactical confirmed kidnook campaign Hudson stapuffer reinforcements Paris Concentraro’s theater! stimket made water excavation blokers Estate Vector Vancouver British infantry company merchant banker subsidiary amended LNG Ferdinand mates uber Schaap m Royalty fracture PSA Conv drafts navigate Parse Site-name CrossRef SC__\__K1-apemiller__\__MP Ref lightweight winds Hurricane winds login Joint GetString Parameters disparities Orth Rocket Venting MPI resemble are Met Lev arc-str sand erosion culernels Hophobic Inbox ashes Cosmos shaping Open whitespace subsidizing Urprot Monthly-Stagg NZ archivetiles coastline-connected Stretch Tribunal Recent Signing exposing Directors rose reveal FA corp Sew pro Last ranks banned Tokibi FusionRib bath storageSettings metaValidateFallback macros Un subtitle Rut Mexican commentary Ribad uploading grow encryption reading Util classes Teaching Alternative indent workflowsJSON filepath Strings testBy Samplefree textile Parser elem pract OakWhen nodes Up representatives Knoxville ODEM repositories BP fixed role Renighbours EIF Recall Copy Destruction gears
3
The instruction was to determine the correct output based on a given input and its corresponding choices.
The instruction was to determine the output choice based on the given inputs and options.
The instruction was to select the choice with the highest point value.
The instruction was to select one choice from each list and provide the selected choice as the output.
The instruction was to select the correct choice from each input sequence.
4
RSelect the correct output film title from the given list of input films.
Choose the correct title from a list of options.
5 Select one movie from the given choices based on the input movies.
Determine the correct movie choice based on the given options for each input.
Given a list of movie titles, you need to choose the correct movie from the given choices that matches closely with the given titles.
Find the correct output movie from the provided choices.
Select the movie from the given choices.
Table 4: Clusters for “rhymes”: the best prompt overall is marked in red, and the best prompt in each cluster in yellow.
Cluster
Prompts
0
The instruction was to identify any homophones in the given inputs.
Identify the words from the given inputs.
1
The instruction was to find the nearest rhyming word for each given word.
Find the rhyming word for the given word.
Replace the provided word with a similar word that rhymes with it and has a different meaning.
Find the words that rhyme with the given word.
The instruction was to provide the output word that rhymes with the given input word.
2 The instruction was to identify any words that are still the same after removing the letters ”in”, ”re”, ”pro”, ”ex”, and ”anti” from the beginning or middle of the word.
Identify the words that are pronounced the same but have different meanings (homophones).
Identify the incorrect word within each pair.
Identify any words that are one letter away from a real word that makes sense.
Identify the word that can be created by changing a single letter at a time from the given word.
3
Find the rhyming word for each input word.
The instruction was to find the rhyming word for each input word.
Find the rhyming word for each input Identify the rhyming word for each given input word.
Find rhyming words for the given inputs.
Find the rhyming word for each input word.
Generate rhyming words with the given input words.
Find the rhyming word for each input word.
4
Replace the word ’phone’ with a similar word.
Identify the words that rhyme with ”phone”.
Identify the words that rhyme with ”phone”.
Identify words that rhyme with ’phone’ and suggest the alternative word that rhymes with each inputted word.
The instruction was to list the words that rhyme with ”phone”.
5
Provide the correct spelling for the given words.
Correct the spelling of the word if it is misspelled, otherwise, keep the word as it is.
Identify the correct spelling of the word.
Replace the letter ”o” with the letter ”a” in each word.
The instruction was to correct any misspelled words in the given inputs.
Figure 2: Complete results with small candidate pools $|{\mathcal{P}}|=10$ and gpt-3.5 on the Instruction-Induction dataset under a budget of $50$.
Figure 3: Complete results with large candidate pools $|{\mathcal{P}}|=30$ and gpt-3.5 on the Instruction-Induction dataset under a budget of $150$.
Figure 4: Complete results with large candidate pools $|{\mathcal{P}}|=30$ and llama-7b on the Instruction-Induction dataset under a budget of $150$.
Figure 5: Complete results with small candidate pools $|{\mathcal{P}}|=10$ and gpt-3.5 on the BigBench-ii dataset under a budget of $50$.
Figure 6: Complete results with large candidate pools $|{\mathcal{P}}|=10$ and gpt-3.5 on the BigBench-ii dataset under a budget of $150$.
Figure 7: Complete results with large candidate pools $|{\mathcal{P}}|=30$ and llama-7b on the BigBench-ii dataset under a budget of $150$.
Table 5: The ratios of different methods outputting a good prompt with gpt-3.5 from large prompt pools $|{\mathcal{P}}|=30$.
[…]
Cause and effect 5 20 20 20 30 60 30
10 20 20 40 40 80 40
20 60 60 80 80 100 80
Common concept 5 0 0 0 0 20 40
10 20 20 20 20 60 40
20 40 40 80 80 80 80
Larger animal 5 80 80 100 100 100 100
10 100 100 100 100 100 100
20 100 100 100 100 100 100
Informal to formal 5 0 0 0 35 25 30
10 20 20 20 60 40 40
20 60 60 80 100 100 100
Negation 5 90 100 100 100 100 100
10 100 100 100 100 100 100
20 100 100 100 100 100 100
Rhymes 5 10 10 30 20 40 30
10 40 40 60 60 80 80
20 100 100 100 100 100 100
Orthography starts with 5 30 40 40 20 40 40
10 60 60 80 80 80 80
20 100 100 100 100 100 100
Sentence similarity 5 25 30 40 25 55 45
10 40 40 60 60 60 80
20 60 60 80 80 80 100
Word in context 5 55 55 70 60 70 70
10 100 100 100 100 100 100
20 100 100 100 100 100 100
Disambiguation qa 5 80 90 100 100 90 100
10 100 100 100 100 100 100
20 100 100 100 100 100 100
Gender Inc. DE 5 40 60 70 80 100 80
10 60 80 100 100 100 100
20 100 100 100 100 100 100
Hyperbaton 5 65 70 70 70 90 80
10 80 80 80 80 100 100
20 100 100 100 100 100 100
Movie recommendation 5 20 20 25 45 50 40
10 40 40 40 60 60 60
20 60 60 80 80 80 80
Object counting 5 10 20 25 30 35 35
10 20 40 40 60 60 60
20 80 100 100 100 100 100
Question selection 5 0 0 10 0 20 15
10 20 20 20 20 40 20
20 40 40 40 60 80 60
Snarks 5 0 20 10 25 25 20
10 20 40 40 60 60 60
20 80 100 100 100 100 100
Word sorting 5 55 70 80 80 75 80
10 100 100 100 100 100 100
20 100 100 100 100 100 100
Ruin names 5 35 55 70 75 70 80
10 60 80 100 100 100 100
20 100 100 100 100 100 100
Sporting understanding 5 75 80 80 80 80 90
10 100 100 100 100 100 100
20 100 100 100 100 100 100
Word unscrambling 5 80 85 90 90 85 90
10 100 100 100 100 100 100
20 100 100 100 100 100 100